OpenStudy (anonymous):

need help with algebra x-8/x+3=8/9

OpenStudy (anonymous):

how much can you do?

hero (hero):

Hint: Cross multiply

OpenStudy (anonymous):

@jmhan222 \[\frac{ x-8 }{ x+3 } = \frac{ 8 }{ 9 }\]

OpenStudy (anonymous):

so you could cross multiply 9 (x-8) = 8 (x + 3) ..then simplify it, and solve for x.

OpenStudy (anonymous):

thanks can you break it down step by step i am having a hard time